# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v2

@@ V-A5 Sensitivity to misalignment of the reference
Figure 10: Sample slice from the LUMC-COR recon-test set comparing PnP-MUNIT with PnP-CNN and MUNIT image translation.
V-A6 Tuning the CR step size
The hyperparameter $\gamma$ in Algorithm 1 controls the CR strength. Fig. 6 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal $\gamma$ decreased with the amount and quality of k-space data. This makes sense as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
V-B Benchmark on NYU DICOM Data
Fig. 7 shows evaluation metrics for PnP-MUNIT and the end-to-end baselines on the NYU DICOM dataset. Despite being trained only on image-domain data, PnP-MUNIT largely outperformed MoDL ($p<0.05$ throughout, except SSIM at $R$=2 and DISTS at $R\in${2,3,4}) and MTrans ($p<$0.05), and was roughly comparable to MC-VarNet. Moreover, as shown in Fig. 8a, PnP-MUNIT could resolve certain fine details even at the high acceleration of $R$=5, which the baselines failed at. In the pathological case shown in Fig. 8b, PnP-MUNIT produced sharper reconstructions preserving the structure of the edema better than MC-VarNet despite lower metrics when many features of edema were absent in the reference image.
V-C Benchmark and Ablation on LUMC Multi-coil Data
@@
The LUMC datasets represent a more challenging problem with real-world data constraints. Additionally, in the light of the content resolution analysis (Section V-A2), we empirically found the ContentRes-2 model configuration as optimal for the LUMC datasets, compared to NYU DICOM data for which ContentRes-1 was the optimal model configuration. This reflects the lower effective content in LUMC image data and hence a greater difficulty for guided reconstruction.
Fig. 9 plots the benchmark metrics for the LUMC test sets, which can be summarized in the following three trends. First, pure image translation with MUNIT was worse compared to single-contrast reconstruction, especially at lower acceleration factors (comparing with PnP-CNN, $p$$<$0.05 for all metrics and both datasets). Combining it with L2-regularized reconstruction in PROSIT improved SSIM and HaarPSI ($p$$<$0.05 for both datasets and accelerations), but not necessarily DISTS, suggesting that the the available complementary information was not fully utilized. PnP-MUNIT was consistently better than both MUNIT image translation and PROSIT ($p$$<$0.05 throughout for both cases). It also outperformed the conventional guided CS-STV ($p$$<$0.05 throughout, except SSIM at $R$=4 in LUMC-COR). Hence, using both measured k-space and reference scan via PnP-MUNIT was more beneficial than (a) using either one of them and (b) combining both using hand-crafted priors, suggesting that our approach maximally exploits the complementary information. Second, both PnP-CNN and PnP-MUNIT performed similarly at lower acceleration, except in terms of SSIM and HaarPSI at $R$=4 and $R$=6 on LUMC-COR where PnP-CNN was slightly better. At higher acceleration, PnP-MUNIT outperformed PnP-CNN ($p$$<$0.05 for all metrics and both datasets except SSIM at $R$=8 in LUMC-COR). Compared to PnP-CNN, PnP-MUNIT allowed up to 32.6% more acceleration for a given SSIM. [^3] Based on linear interpolation of LUMC-TRA SSIM, PnP-MUNIT and PnP-CNN allowed $R$=10 and $R$=6.7, the maximum difference in $R$, at median SSIM=0.906. Hence, a difference of 32.6%. Third, introducing PFT and CR into the ablated PnP-MUNIT improved the reconstructions ($p$$<$0.05 throughout in both cases except in DISTS for the latter at $R$=10 in LUMC-TRA). The contribution of PFT was almost constant across $R$, whereas that of CR decreased with $R$, which was expected since CR depends on the measured k-space data to refine the content.
Fig. 10 shows an LUMC-COR example representative of the first two trends. MUNIT image translation produced severe anatomical defects and predicted false structure in the lower region of the image where the ground truth, in fact, contains low signal. At $R$=4, PnP-MUNIT was visually similar to PnP-CNN, with the main difference being a blur effect in PnP-CNN and mild texture artifacts in PnP-MUNIT. At $R$=10, PnP-CNN reconstruction was non-viable, containing severe blur and artifacts (green arrows). On the other hand, PnP-MUNIT did not degrade much from the $R$=4 case and most fine structures were sharply resolved (green arrows). However, an interesting failure mode of PnP-MUNIT at high acceleration was the subtle localized distortions in the anatomy (pink arrows).
@@
Figure 11: Averaged results of the radiological evaluation.
V-D Radiological Evaluation
@@
Figure 11 plots the radiological evaluation result. At $R$=3, both algorithms produced at least “fairly diagnostic” reconstructions and allowed the detection of more than 90% of metastases, a difference of 33.3% of k-space samples over the clinical reconstructions of $R$=2. PnP-CNN was better than PnP-MUNIT at $R\in\{3,4\}$ in terms of sharpness, contrast, and the pathology criteria. At $R$=5, PnP-MUNIT matched or exceeded PnP-CNN in terms of visual quality, although being worse in terms of pathology criteria. At $R$=6, PnP-CNN sharply dropped to non-diagnostic level, whereas PnP-MUNIT was better, especially in terms of sharpness and contrast as well as in overall diagnostic quality.
VI Discussion
@@
In this work, we modeled the two-contrast MR image data in terms of two latent generative factors – content representing the contrast-independent structure and style representing the contrast-specific variations. Our data-driven definitions of content and style, though seemingly related to the MR physics-based representations of quantitative maps and acquisition-related factors, are rather nebulous compared to precisely defined physical concepts. For instance, MR quantitative maps are theoretically objective representations of tissues in terms of physical variables such as relaxation times and are ideally independent of the contrast-generating sequence parameters. On the other hand, our content and style representations are only defined statistically in terms of the given two contrasts and are, thus, neither as objective nor as independent in the ideal physical sense. That being said, we would argue that there is potential value in augmenting the data-driven content/style model with MR physical models, e.g. by constraining the learned content to represent physically meaningful anatomical properties, introducing elements of a physical model (e.g. Bloch equation) into the decoder network, etc., to enhance interpretability and reconstruction quality.
@@
Optimal content capacity was defined as a quantity representing the amount of shared local information contained in a two-contrast dataset. In terms of this quantity, we analyzed the effect of reference image resolution on reconstruction quality, obtaining a lowerbound for PnP-MUNIT to be effective (Section V-A2). This quantity would depend on more fundamental factors such as MR sequence types, which we did not investigate here. While we limited our experiments to T1W and T2W sequences, PnP-MUNIT is applicable to any pair of contrasts, with the image quality limited mainly by the amount of structural information shared between them. Future work could help empirically determine which sequence pairs are more amenable to guided reconstruction than others. Note that for a different pair of contrasts, say PD-weighted and T2W, a fresh content/style model needs to be trained since content and style are defined in the context of the specific contrast pair. It is, in principle, also possible to model the content and style for more than two contrasts. Given number of contrasts $N$, content is defined as the local structure underlying all $N$ contrasts, while style encodes intra-contrast variations of each individual contrast. This model would include $N$ sets of encoders and decoders and would require an (unpaired) image-domain training dataset of these $N$ contrasts. Subsequently, our PnP-MUNIT framework can be extended to use multiple reference contrasts to guide a given target contrast. The specific question to investigate then would be – given $P$ reference contrasts (where $1<P<N$), how to aggregate their contents such that the prior structural information about the target image is maximized?
@@
Content discrepancy was characterized as a quantity representing the gap between the true content of the target image and our estimation of this content, which is another limiting factor in PnP-MUNIT reconstruction. In MUNIT, the decoders model the (forward) generative process that maps the latent variables (i.e. content and style) to the observable variable (i.e. images). Ideally, the encoders must be a perfect inverse of this forward process. However, estimating the multi-channel contrast-independent content from a single image (using the content encoders) is a challenging (and perhaps ill-conditioned) inverse problem. Hence, some errors are to be expected in the content estimated from the reference image. Moreover, the reference image may contain unfavorable differences from the target image, such as misalignment or artifacts, which introduce more errors. Therefore, content discrepancy is expected to be a non-zero quantity. We proposed two complementary ways to minimize its effect on the reconstruction. While paired fine-tuning improves the model using a small amount of paired image-domain training data, content refinement (CR) aims to correct the remaining discrepancy during reconstruction using the measured k-space. A drawback of the CR module, however, is slower convergence (Section V-A3). For example, the run-time for PnP-MUNIT on the GPU was around 17 seconds for a slice of matrix size 349$\times$284, as compared to 6 seconds for CS-WT, 1.4 seconds for PnP-CNN, and less than 0.1 second for MC-VarNet. This latency issue could be mitigated by approximating the iterative process with an unrolled network design, although perhaps at a cost of the generalizability offered by the plug-and-play design. The resulting trade-off between model latency and generalizability would be interesting to explore in the future.
@@
In terms of training data requirements, PnP-MUNIT can make use of the larger amounts of the image data available and can be applied to situations where end-to-end methods are infeasible since it does not rely on k-space training data. Moreover, unpaired image-domain pre-training of the content/style model boosts the practical applicability of PnP-MUNIT, e.g. in the case of LUMC-TRA where the T1W/T2W data imbalance was considerable. As indicated in our NYU DICOM benchmark, PnP-MUNIT performs similarly or better than state-of-the-art end-to-end methods. In our LUMC multi-coil benchmark, we limited our baselines to those algorithms that were feasible given the data constraint. As future work, a more comprehensive study should be conducted with a broader range of methods and evaluation criteria, e.g. comparing generalizability across accelerations and sampling patterns with multi-contrast unrolled networks, given a fixed budget of paired training data. Contrary to the training stage, we strictly assumed the spatial alignment of the reference and target images in the reconstruction stage. The CR update, which can implicitly correct minor registration errors, would likely break down at the typically observed levels of patient motion. A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further improving practical applicability.
@@
In terms of reconstruction quality, we observed in Section V-C that the true added value of PnP-MUNIT was at the high acceleration factors ($R$=8 and $R$=10) where the k-space data is scarce and the reference information becomes more valuable. However, the risk of model hallucinations increases at these accelerations, raising concerns about the accuracy of the anatomy represented in the image. Given the advantages of content/style decomposition, it may be possible to leverage the contrast-independence property of the content to automatically detect local hallucinations in the reconstruction and to define a corrective process to minimize it. This is another topic for future work. At lower accelerations, on the other hand, we observed that PnP-MUNIT often produced slightly lower metrics compared to PnP-CNN, e.g. in LUMC-COR. This was also observed in the radiological evaluation on LUMC-TRA samples. This is counter-intuitive at first glance since PnP-MUNIT has access to additional side-information and should, in principle, perform at least as good as PnP-CNN. However, this can be explained by the fact that at lower accelerations, the content/style model becomes the limiting factor for reconstruction quality. This is because unlike the CNN denoiser (Eq. (4)) of PnP-CNN, the content consistency operator (Eq. (14)) of PnP-MUNIT radically changes the image and hence, this operation is sensitive to the model’s overall performance. The model performance is in turn influenced by the training dataset and the network architecture. First, regarding data, the T1W scans in LUMC datasets had an in-plane resolution 1.3-2.5 times lower than that of the T2W scans, limiting the quality of the model (as indicated by its low optimal content capacity) and the overall value of guidance. In the long run, this may be solved by using 3D reference and target sequences of matched-resolution to optimize guidance. Second, regarding model design, correcting for content discrepancy using the CR module relies on computing gradients through the decoder $G_{2}$ (Eq. (16)). If $G_{2}$ is highly non-linear and the reference content is not sufficiently close to the optimum, CR may converge to a sub-optimal point leading to texture artifacts in the reconstruction, as seen in Fig. 10 at $R$=4. In our proof-of-concept, we used the same general model architecture proposed in [22]. Better designs that are more conducive to CR may exist. Hence, future efforts must considerably focus on improving the content/style model using specialized architecture and regularization.
VII Conclusion
@@
In this work, we introduced PnP-MUNIT, a modular approach to multi-contrast reconstruction combining content/style modeling with iterative reconstruction, which offers the reconstruction quality of end-to-end methods under stronger training data constraints. At its core is the content consistency operation, which provides regularization at the level of the image’s semantic content. We defined two quantities that determine the efficacy of this operation, namely optimal content capacity and content discrepancy, and provided several ways of maximizing this efficacy. On real-world clinical data, PnP-MUNIT provided up to 32.6% more acceleration over PnP-CNN for given SSIM, enabling sharper reconstructions at high accelerations. In the radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-MUNIT produced diagnostic-quality images at $R$=3, enabling 33.3% more acceleration over clinical reconstructions. To progress towards practical implementation of our proof-of-concept, future work will focus mainly on improving the content/style model and reconstruction latency, tackling the problem of model hallucinations, and incorporating online registration into the reconstruction.

### v3

@@ 5.3 Benchmark and ablation on the LUMC multi-coil data
Figure 11: Sample slice from the LUMC-COR recon-test set comparing PnP-CoSMo with PnP-CNN and MUNIT image translation.
@@
The LUMC datasets represent a more challenging problem with real-world data constraints. Additionally, in the light of the content capacity analysis (Section 5.1.2), we empirically found that the content capacity $J_{M}$=1/4 was optimal for the LUMC datasets, compared to NYU DICOM data where $J_{M}$=1 was the optimal content capacity. This reflects the lower effective content in LUMC image data and hence a greater difficulty of the guided reconstruction problem.
@@
Figure 12: Scores from the radiological evaluation averaged over the three subjects.
Figure 13: Example slice from the LUMC-TRA subset used in the radiological evaluation. Several relevant features are marked including a brain metastasis-related hyperintense feature (salmon-pink arrows), reconstruction artifacts (cyan arrows), and blurring (purple arrows).
Fig. 10 plots the benchmark metrics for the LUMC test sets, which can be summarized in the following three trends. First, pure image translation with MUNIT was worse compared to single-contrast reconstruction, especially at lower acceleration factors (comparing with PnP-CNN, $p$$<$0.05 for all metrics and both datasets). Combining it with L2-regularized reconstruction in PROSIT improved SSIM and HaarPSI ($p$$<$0.05 for both datasets and accelerations), but not necessarily DISTS, suggesting that the the available complementary information was not fully utilized. PnP-CoSMo was consistently better than both MUNIT image translation and PROSIT ($p$$<$0.05 throughout for both cases). It also outperformed the conventional guided CS-STV ($p$$<$0.05 throughout, except SSIM at $R$=4 in LUMC-COR). Hence, using both measured k-space and reference scan via PnP-CoSMo was more beneficial than (a) using either one of them and (b) combining both using hand-crafted priors, suggesting that our approach maximally exploits the complementary information. Second, both PnP-CNN and PnP-CoSMo performed similarly at lower acceleration, except in terms of SSIM and HaarPSI at $R$=4 and $R$=6 on LUMC-COR where PnP-CNN was slightly better. At higher acceleration, PnP-CoSMo outperformed PnP-CNN ($p$$<$0.05 for all metrics and both datasets except SSIM at $R$=8 in LUMC-COR). Compared to PnP-CNN, PnP-CoSMo allowed up to 32.6% more acceleration for a given SSIM. [^2] Based on linear interpolation of LUMC-TRA SSIM, PnP-CoSMo and PnP-CNN allowed $R$=10 and $R$=6.7, which is the maximum difference in $R$, at median SSIM=0.906. Hence, a difference of 32.6%. Third, introducing PFT and CR into the ablated PnP-CoSMo improved the reconstructions ($p$$<$0.05 throughout in both cases except in DISTS for the latter at $R$=10 in LUMC-TRA). The contribution of PFT was almost constant across $R$, whereas that of CR decreased as $R$ increased, which was expected since CR depends on the measured k-space data to refine the content.
Fig. 11 shows an LUMC-COR example representative of the first two trends. MUNIT image translation produced severe anatomical defects and predicted false structure in the lower region of the image where the ground truth, in fact, contains low signal. At $R$=4, PnP-CoSMo was visually similar to PnP-CNN, with the main difference being a blur effect in PnP-CNN and mild texture artifacts in PnP-CoSMo. At $R$=10, PnP-CNN reconstruction was non-viable, containing severe blur and artifacts (green arrows). On the other hand, PnP-CoSMo did not degrade much from the $R$=4 case and most fine structures were sharply resolved (green arrows). However, an interesting failure mode of PnP-CoSMo at high acceleration was the subtle localized distortions in the anatomy (magenta arrows).
5.4 Radiological evaluation
@@
Figure 12 plots the radiological evaluation result. At $R$=3, both algorithms produced diagnostic-quality reconstructions, a difference of 33.3% of k-space samples over the clinical reconstructions of $R$=2. PnP-CNN was slightly better than PnP-CoSMo at $R\in\{3,4\}$ in terms of sharpness, contrast, and the pathology criteria. At $R$=5, PnP-CoSMo matched or exceeded PnP-CNN in terms of visual quality, although being worse in terms of pathology criteria. At $R$=6, PnP-CNN sharply dropped to the ”non-diagnostic” level, whereas PnP-CoSMo was better, especially in terms of sharpness and contrast as well as in overall diagnostic quality. This trend is illustrated with a representative example in Figure 13. While the $R$=3 reconstructions of PnP-CNN and PnP-CoSMo are of a similarly high quality, reconstruction artifacts (cyan arrows) become more pronounced in PnP-CNN reconstructions with increasing acceleration. At $R$=6, compared to PnP-CoSMo, the PnP-CNN reconstruction has severe artifacts, blurrier and more distorted features (purple arrows), and a poorly visible hyperintense area of a brain metastasis (salmon-pink arrows).
6 Discussion
@@
In this work, we modeled two-contrast MR image data in terms of two latent generative factors – content representing the contrast-independent structure and style representing the contrast-specific variations. Our data-driven definitions of content and style, though seemingly related to the MR physics-based representations of quantitative maps and acquisition-related factors, are rather nebulous compared to precisely defined physical concepts. For instance, MR quantitative maps are theoretically objective representations of tissues in terms of physical variables such as relaxation times and are ideally independent of the contrast-generating sequence parameters. On the other hand, our content and style representations are only defined statistically in terms of the given two contrasts and are, thus, neither as objective nor as independent in the ideal physical sense. That being said, we would argue that there is potential value in augmenting the data-driven content/style model with MR physical models, e.g. by constraining the learned content to represent physically meaningful anatomical properties, introducing elements of a physical model (e.g. Bloch equations) into the decoder network, etc., to enhance interpretability and reconstruction quality.
@@
Optimal content capacity was defined as a quantity representing the amount of shared local information contained in a two-contrast dataset. In terms of this quantity, we analyzed the effect of reference image resolution on reconstruction quality, obtaining a lowerbound for PnP-CoSMo to be effective (Section 5.1.2). This quantity would depend on more fundamental factors such as MR sequence types, which we did not investigate here. While we limited our experiments to T1W and T2W sequences, PnP-CoSMo is, in principle, applicable to any pair of contrasts, with the image quality limited mainly by the amount of structural information shared between them. Future work could help empirically determine which sequence pairs are more amenable to guided reconstruction than others. A consequence of our MUNIT-based model design was the model’s lack of dependence on the uni-directionality of the guided reconstruction problem, which allowed greater generalizability by design (Section 5.2). It is, in principle, also possible to model the content and style for more than two contrasts. This would enable not only the generalizability of our approach to any combinatorial set of guided reconstruction tasks, but also better estimation of the target contrast’s content by utilizing multiple reference contrasts in the reconstruction. Given number of contrasts $N$, content could be defined as the local structure underlying all these contrasts, while style would encode intra-contrast variations of each individual contrast. This model would then require an (unpaired) image-domain training dataset of the $N$ contrasts. An open question is how to effectively scale the model and training complexity to $N$ contrasts. The PnP-CoSMo algorithm can subsequently be extended to utilize multiple reference contrasts to guide a given target contrast. The specific question to investigate then would be – given $P$ reference contrasts (where $1<P<N$), how to aggregate their contents such that the prior structural information about the target image is maximized?
@@
Content discrepancy was characterized as a quantity representing the gap between the true content of the target image and our estimation of this content, which is another limiting factor in PnP-CoSMo reconstruction. In our MUNIT-based network design, the decoders model the (forward) generative process that maps the latent variables (i.e. content and style) to the observable variable (i.e. images). Ideally, the encoders must be a perfect inverse of this forward process. However, estimating the multi-channel contrast-independent content from a single image (using the content encoders) is a challenging (and perhaps ill-conditioned) inverse problem. Hence, some errors are to be expected in the content estimated from the reference image. Moreover, the reference image may contain unfavorable differences from the target image, such as misalignment or artifacts, which introduce more errors. Therefore, content discrepancy is expected to be a non-zero quantity. We proposed two complementary ways to minimize its effect on the reconstruction. While paired fine-tuning improves the model using a small amount of paired image-domain training data, content refinement (CR) aims to correct the remaining discrepancy during reconstruction using the measured k-space. A drawback of the CR module, however, is slower convergence (Section 5.1.3). For example, the run-time for PnP-CoSMo on the GPU was around 17 seconds for a slice of matrix size 349$\times$284, as compared to 6 seconds for CS-WT, 1.4 seconds for PnP-CNN, and less than 0.1 second for MC-VarNet. This latency issue could be mitigated by approximating the iterative process with an unrolled network design, although potentially at a cost of the generalizability offered by the plug-and-play design. The resulting trade-off between model latency and generalizability would be interesting to explore in the future.
@@
In terms of training data requirements, PnP-CoSMo can make use of the larger amounts of the image data available and can be applied to situations where end-to-end methods are infeasible since it does not rely on k-space training data. In situations where end-to-end methods are indeed feasible, PnP-CoSMo is comparable to state-of-the-art methods in terms of metrics and superior in terms of visual quality, especially at higher accelerations (Section 5.2 ). Moreover, unpaired image-domain pre-training of the content/style model boosts the practical applicability of PnP-CoSMo, e.g. in the case of LUMC-TRA where the T1W/T2W data imbalance was considerable. As indicated in our NYU DICOM benchmark, PnP-CoSMo performs similarly or better than state-of-the-art end-to-end methods. In our LUMC multi-coil benchmark, we limited our baselines to those algorithms that were feasible given the data constraint. As future work, a more comprehensive study should be conducted with a broader range of methods and evaluation criteria, e.g. comparing generalizability across accelerations and sampling patterns with multi-contrast unrolled networks, given a fixed budget of paired training data. Contrary to the training stage, we strictly assumed the spatial alignment of the reference and target images in the reconstruction stage. The CR module, which can implicitly correct minor registration errors, would likely break down at the typically observed levels of patient motion. A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further improving practical applicability.
@@
In terms of reconstruction quality, we observed in Section 5.3 that the true added value of PnP-CoSMo was at the high acceleration factors ($R$=8 and $R$=10) where the k-space data is scarce and the reference information becomes more valuable. However, the risk of model hallucinations increases at these accelerations, raising concerns about the accuracy of the anatomy represented in the image. We hypothesize that as a consequence of content/style decomposition, it may be possible to leverage the contrast-independence property of the content to – (a) automatically detect local hallucinations in the reconstruction and (b) define a corrective process in combination with an adaptive sampling strategy to minimize it. This is yet another topic for future work. At lower accelerations, on the other hand, we observed that PnP-CoSMo often produced slightly lower metrics compared to PnP-CNN, e.g. in LUMC-COR. This was also observed in the radiological evaluation on LUMC-TRA samples. This is counter-intuitive at first glance since PnP-CoSMo has access to additional side information and should, in principle, perform at least as well as PnP-CNN. However, this can be explained by the fact that at lower accelerations, the content/style model becomes the main limiting factor for reconstruction quality. This is because unlike the CNN denoiser (Eq. (4)) of PnP-CNN, the content consistency operator (Eq. (15)) of PnP-CoSMo radically alters the image and hence, this operation is sensitive to the model’s overall performance. The model performance is in turn influenced by the training dataset and the network architecture. First, regarding data, the T1W scans in LUMC datasets had an in-plane resolution 1.3-2.5 times lower than that of the T2W scans, limiting the quality of the model (as indicated by its low optimal content capacity) and the overall value of guidance. In the long run, this may be solved by using 3D reference and target sequences of matched-resolution to optimize guidance. Second, regarding model design, correcting for content discrepancy using the CR module relies on computing gradients through the decoder $G_{2}$ (Eq. (17)). If $G_{2}$ is highly non-linear and the reference content is not sufficiently close to the optimum, CR may converge to a sub-optimal point leading to texture artifacts in the reconstruction, as seen in Fig. 11 at $R$=4. In our proof-of-concept, we used the same general model architecture proposed in Huang et al. (2018). Better designs that are more conducive to CR may exist. Hence, future efforts must considerably focus on improving the content/style model using specialized architecture and regularization.
7 Conclusion
@@
In this work, we introduced PnP-CoSMo, a modular approach to multi-contrast reconstruction which combines content/style modeling with iterative reconstruction, offering the reconstruction quality of state-of-the-art end-to-end methods with improved generalizability despite lenient training data requirements. At its core is the content consistency operation, which provides regularization at the semantic level of the image content. We defined two quantities that determine the efficacy of this operation, namely optimal content capacity and content discrepancy, and provided several ways of maximizing this efficacy. On real-world clinical data, PnP-CoSMo provided up to 32.6% more acceleration over PnP-CNN for given SSIM, enabling sharper reconstructions at high accelerations. In the radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-CoSMo produced diagnostic-quality images at $R$=3, enabling 33.3% more acceleration over clinical reconstructions. To progress towards practical implementation of our proof-of-concept, future work will focus mainly on improving and extending the content/style model, reducing reconstruction latency, tackling the problem of model hallucinations, and incorporating online registration into the reconstruction.
CRediT authorship contribution statement
Chinmay Rao: Conceptualization, Methodology, Data curation, Software, Investigation, Validation, Formal analysis, Visualization, Writing – original draft. Matthias van Osch: Funding acquisition, Supervision, Conceptualization, Validation, Writing – review & editing. Nicola Pezzotti: Funding acquisition, Supervision, Conceptualization, Methodology, Validation. Jeroen de Bresser: Investigation, Validation. Laurens Beljaards: Data curation, Software. Jakob Meineke: Resources, Software. Elwin de Weerdt: Conceptualization, Resources. Huangling Lu: Investigation. Mariya Doneva: Supervision, Conceptualization, Methodology, Validation, Writing – review & editing. Marius Staring: Funding acquisition, Project administration, Supervision, Conceptualization, Methodology, Validation, Writing – review & editing.
Declaration of competing interest
The authors declare the following financial interests or personal relationships which may be considered as potential competing interests: Chinmay Rao and Laurens Beljaards declare the grant and research support they receive received from Philips. Nicola Pezzotti, Jakob Meineke, Elwin de Weerdt, and Mariya Doneva declare their employment at Philips. Given his role as Associate Editor, Marius Staring had no involvement in the peer review of this article and had no access to information regarding its peer review. Full responsibility for the editorial process for this article was delegated to another journal editor. If there are other authors, they declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work is part of the project ROBUST: Trustworthy AI-based Systems for Sustainable Growth with project number KICH3.LTP.20.006, which is (partly) financed by the Dutch Research Council (NWO), Philips Research, and the Dutch Ministry of Economic Affairs and Climate Policy (EZK) under the program LTP KIC 2020-2023.
Data availability
The authors do not have permission to share data.